OpenStudy (anonymous):

PLEASE HELP ME I WILL MEDAL AND FAN!!!!!!!!!!!!!!!!! Mr. Evans paid $455 for a 7-day car rental. At that rate, how much would he pay for a 10-day rental? A.$465 B. $462 C. $65 D. $650 I think its A

OpenStudy (anonymous):

@bohotness

OpenStudy (bohotness):

subtract love

OpenStudy (anonymous):

subtract what

OpenStudy (bohotness):

17 with 445

OpenStudy (anonymous):

445-17=428 @bohotness

OpenStudy (bohotness):

add 445 with 7 subtract 10

OpenStudy (anonymous):

442 honey lol @bohotness

OpenStudy (anonymous):

they are messing with you

OpenStudy (anonymous):

so what do i do??

OpenStudy (anonymous):

\[\frac{455}{7}=\frac{x}{10}\\ x=\frac{10\times 455}{7}\]

OpenStudy (anonymous):

whats the x for??

OpenStudy (anonymous):

the number you are looking for

OpenStudy (anonymous):

$455 for 7 days is \(455\div 7=65\) dollars a day

OpenStudy (anonymous):

can u give me the answer cause i have 2 go like now

OpenStudy (anonymous):

if you rent at $65 per day for 10 days you pay \[10\times 65=650\] dollars